Given the following data, answer the questions below the table: ------------------------------------------------------------------------------------------------------------------------ Data Allocation base Machine-hours Estimated...

  1. Given the following data, answer the questions below the table:

------------------------------------------------------------------------------------------------------------------------

Data

Allocation base

Machine-hours

Estimated manufacturing overhead cost

$220,800

Estimated total amount of the allocation base

46,000

machine-hours

Actual manufacturing overhead cost

$240,100

Actual total amount of the allocation base

42,000

machine-hours

------------------------------------------------------------------------------------------------------------------------

a) What is the pre-determined overhead rate?

b) By how much is the manufacturing overhead underapplied or overapplied?

c) Change the estimated total amount of allocation base to 48,000 machine hours, now by how much is the manufacturing overhead underapplied or overapplied?
Answer #1

Predetermined Overhead rate = Estimated overhead/Estimated machine hours

= 220,800/46,000

= 4.80 per machine hour

Applied overhead = 4.80 * 42,000 = 201,600

Actual overhead = 240,100

So, overhead underapplied by = 240,100 - 201,600

= 38,500 underapplied

Predetermined overhead rate = 220,800/48,000 = 4.60 per mh

Applied overhead = 4.60*42000 = 193,200

Actual overhead = 240,100

So,overhead is underapplied by = 240,100 - 193,200

= 46,900 underapplied